# Temporal reasoning for timeline summarisation in social media

### v1

@@ 4.1 Datasets
We conduct experiments on three different datasets. We fine-tune the teacher model on the ‘NarrativeReason’ dataset. When distilling the temporal reason knowledge to the student model, we use a news timeline summarisation datasetChen et al. (2023a). Finally, we apply the model to a different domain, specifically generating mental health-related timeline summaries from social media.
@@
Timeline summarisation Dataset This timeline summarisation dataset is sourced from Chen et al. (2023a). They collected timeline summaries from Wikipedia websites, with a total of 5,000 timelines and summaries.
TalkLife When generating the summary, we use the dataset collected by Tsakalidis et al. (2022) comprising 500 anonymised user timelines from Talklife. Song et al. (2024) sample 30 timelines from it and annotate them with corresponding mental health-related summaries. These human-written summaries include diagnosis, intra- and interpersonal patterns and mental state changes over time, and they also highlight information related to individuals’ mental states, which can be used for automated evaluation.
@@ 4.2 Models & Baselines
We compare our method against existing LLMs for mental health related summarisation. We introduce the implementation details in the appendix A.1
L-phi This model uses LLaMA as the teacher model and a smaller model, Phi, as the student. We apply different knowledge distillation (KD) methods to transfer temporal reasoning knowledge to the student model. Subsequently, we directly prompt this model to generate mental health-related timeline summaries.
P-phi In this model, we use Phi as the teacher model and another Phi of the same size as the student model.
$\text{Phi}_{joint}$ To compare with the phi-phi, we use joint learning to fine-tune Phi on both the NarrativeReason and Timeline summarisation datasets. This allows us to observe whether knowledge distillation (KD) outperforms directly fine-tuning on the two datasets. In addition, we fine-tune Phi on the NarrativeReason and timeline summarisation datasets and obtain models $\text{Phi}_{temp}$ and $\text{Phi}_{timeline}$ separately. We use these two models to generate timeline summaries for comparison with the KD derived model.
$\text{Phi}_{ICL}$ We use in context learning (ICL) to guide Phi to generate summaries. We provide the model with a pair consisting of a timeline and its corresponding summary as an example, and then let it generate summaries for other timelines.
LLaMA We prompt LLaMA to generate mental health related summaries on timelines directly as a baseline.
@@ 4.3 Evaluation
We use the timeline summaries from Song et al. (2024)4.1) forautomated evaluation. We employFactual Consistency (FC), to measure whether timeline summaries are consistent with the original timelines, and Evidence Appropriateness (EA), to measure the consistency of human written summaries with their corresponding timeline summaries Song et al. (2024).
For human evaluation, we worked with two clinical psychology graduate students fluent in English to evaluate 30 summaries generated from 30 timelines (TalkLfe). We follow the metrics used in Song et al. (2024) to evaluate the summaries from the perspectives of Factual Consistency and Usefulness (general/diagnosis/inter-&Interpersonal/MOC). A factually consistent summary should accurately represent the content of the timeline and excludes any information not present in the timeline. We also evaluate the usefulness of a mental health summary on the basis of four aspects: General (contains the most clinically important information from the timeline and aids the clinician in understanding a patient’s condition); Diagnosis (provides useful information about an individual’s diagnosis); inter-&Interpersonal (provides helpful information about an individuals’ main needs and patterns of self and other relationships); MOC (provides useful information about an individual’s changes over time in terms of emotion/cognition and behaviour).
@@ 5 Results and Discussion
5.1 Automatic evaluation
We conducted experiments with different combinations of KD methods. Table 1 shows experiment results. Since we did not change the output dimensions when fine-tuning LLaMA, the CRD method was not used during the KD process. Among the individual methods, PKT performed the best; however, combining PKT with NST achieved the best overall results.
@@ 5.1 Automatic evaluation
Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) on student model.
Higher is better, best in bold
Table 2 shows the results of applying alternative fine-tuning strategies on LLMs. The results for $\text{Phi}_{temp}$ and $\text{Phi}_{timeline}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues. Notably, $\text{Phi}_{temp}$ performed the worst in both FC and EA metrics, suggesting that the incorporation of temporal reasoning information appears to interfere with the LLM’s ability to effectively handle the timeline summarization task. Additionally, in strategy $\text{Phi}_{jont}$, combining the two types of data directly during training failed to integrate them effectively. As a result, the performance of $\text{Phi}_{jont}$ was even worse than using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
Metric $\text{Phi}_{ICL}$ $\text{Phi}_{temp}$ $\text{Phi}_{timeline}$ $\text{Phi}_{joint}$ FC .412 .141 .184 .238 EA .965 .895 .966 .941
Table 2: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) across the compared models. Higher is better, best in bold
5.2 Human evaluation
Based on the results of the automatic evaluation, we selected the best-performing LLaMA-Phi and Phi-Phi models. Additionally, we included the non-fine-tuned versions of LLaMA and Phi. This can help us understand in which specific aspects the model has improved with the inclusion of temporal reasoning information. The fine-tuned model shows the greatest improvement in terms of factual consistency and usefulness (general). This aligns with our findings when analyzing the summaries, where the fine-tuned model significantly reduces hallucination, as shown in Table 3. In addition, we found that the fine-tuned model did not show significant improvement in terms of Moments of Change (MOC). This could be because, in the timelines, users do not experience many emotional switches, resulting in the model-generated summaries being relatively similar.
Aspect phi phi-phi LLaMA LLaMA-phi Factual Consistency 2.90 3.32 3.58 3.83 Usefulness (General) 2.60 3.13 3.17 3.48    (Diagnosis) 2.90 3.37 3.45 3.62   —-(Inter-& Intrapersonal) 2.95 3.00 3.40 3.51   —-(MoC) 2.97 2.97 3.42 3.47
@@ 5.3 Why knowledge distillation works
In this section, we analyze why the $\text{Phi}_{timeline}$ model performed better from a representation learning perspective. For this purpose, we perform two experiments, a task understanding probing experiment and a Joint Task Representation Learning (JTRL) experiment. We analyze the internal representations of $\text{Phi}_{timeline}$ against $\text{Phi}_{joint}$. We construct our probing dataset by using $\text{Phi}_{joint}$’s test dataset. We pose the two tasks as a binary classification problem and then extract the activations for the last layer. Then we use UMAP (McInnes et al., 2018) to project these activations to lower dimensions (Sainburg et al., 2021; Tseriotou et al., 2023). From Figure 3, we can see that the activations of $\text{Phi}_{joint}$ are well separated for each task, whereas the activations for $\text{Phi}_{timeline}$ overlap. Given the performance of $\text{Phi}_{timeline}$, this leads us to hypothesize that the model learned better representations for the task due to more task specific polysemantic (Olah et al., 2020) neurons.
To validate our hypothesis, we ran another set of experiments to analyze the internal representation difference between the two models, which we termed as JTRL. Our main idea here was to validate that knowledge distillation resulted in better representations due to more polysemantic neurons, and these representations varied highly compared to $\text{Phi}_{joint}$’s representation. To measure JTRL, we use the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) similarity score. CKA can be used to measure the similarity between internal representations of modelsConneau et al. (2020); Muller et al. (2021); Del and Fishel (2021); Moosa et al. (2023).
To calculate CKA, we used the same probing dataset. First, we calculate the sentence embeddings of each input by averaging the hidden state representation of the tokens. Then, we calculate the CKA similarity score between the mean sentence embeddings and each layer representation of the model. We did this both for the individual models and between the layers of the models. From Figure 4, we can see that $\text{Phi}_{timeline}$ shows a gradual increase in CKA across layers, peaking in the mid-to-late layers. This indicates that as the layers progress, $\text{Phi}_{timeline}$ preserves and refines the information from the initial embeddings in a task-relevant way. The is likely due to the fact that the distillation process encourages this alignment by transferring task-relevant knowledge from the teacher. On the other hand, $\text{Phi}_{joint}$ shows an initial increase but saturates and flattens early. This indicates that it fails to refine representations effectively in deeper layers, which could be due to conflicting objectives between the tasks. The lower CKA in later layers suggests that the model moves away from the initial embeddings in a manner that is less effective for task-specific learning. Lastly, the CKA values between the models show that they learn vastly different representations. In short, $\text{Phi}_{timeline}$’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
@@
(b) UMAP projection $\text{Phi}_{timeline}$
@@
Figure 4: CKA similarity score of both within and between $\text{Phi}_{timeline}$ and $\text{Phi}_{joint}$ model representations.
@@ 6 Conclusions
We created a dataset named NarrativeReason, which is utilised to enhance the temporal reasoning abilities of LLMs. This dataset enables models to uncover temporal relationships between events within timelines. We fine-tune a larger LLM with NarrativeReason, then distill its enhanced temporal reasoning capability to a smaller LLM while leveraging the distilled knowledge to enhance performance in timeline summarisation tasks. We apply the model to a different domain from the one it was trained on, namely generating mental health related timeline summaries. Our results demonstrate that our approach using knowledge distillation produces more accurate summaries while significantly reducing hallucinations. We further analyzed the internal representations of the model and found that knowledge distillation leads to better feature representations within the model that are more aligned with the task of timeline summarisation.
@@ A.2 Experiment on teacher model
After fine-tuning the teacher model, we used the TEMPREASON dataset Tan et al. (2023b) to evaluate its performance pre- and post-fine-tuning. Specifically, we focused on the L3 part of the dataset, which deals with event-event relations, to determine whether the model could accurately infer the temporal sequence between two events. We use F1 as the evaluation metrics, the experimental results show that the fine-tuned model achieved a 0.07 improvement in this metric compared to the pre-fine-tuning model, highlighting its enhanced ability to infer event-event temporal relations.

### v2

@@ 4.1 Datasets
We conduct experiments on three different datasets. We fine-tune the teacher model on the ‘NarrativeReason’ dataset. When distilling the temporal reasoning knowledge to the student model, we also fine-tune the latter on a news timeline summarisation dataset Chen et al. (2023). Finally, we apply the fine-tuned student model to a different domain, specifically that of generating mental health-related timeline summaries from social media.
@@
Timeline summarisation Dataset The timeline summarisation dataset used for training comes from Chen et al. (2023). It consists of timeline summaries from Wikipedia websites, with a total of 5,000 timelines and summaries. Due to some inconsistencies between events across timelines and summaries, this dataset was only used for training.
TalkLife When generating the summary, we use the dataset collected by Tsakalidis et al. (2022) comprising 500 anonymised user timelines from Talklife  [^1] https://www.talklife.com/. Song et al. (2024) had sampled 30 timelines from it and augmented them with corresponding mental health-related summaries and associated evidence. The summaries cover aspects such as diagnosis, intra- and interpersonal patterns and mental state changes over time. They also highlighted associated evidence in the timelines, which they utilised for automated summary evaluation.
@@ 4.2 Models & Baselines
For mental health related summarisation, we compare our method against existing LLMs . Implementation details are in appendix A.1
L-Phi: This model derives from a LLaMA-3 teacher model teaching a smaller student model, Phi. We apply different knowledge distillation (KD) methods to transfer temporal reasoning knowledge to the student model. L-Phi is also fine-tuned for timeline summarisation. Subsequently, we directly prompt this model to generate mental health-related timeline summaries.
P-Phi: The same as L-Phi but we use another Phi as the teacher model instead of LLaMA-3.
$\text{Phi}_{joint}$: To investigate the effect of knowledge distillation (KD), we compare P-Phi, against a joint learning setting where we fine-tune Phi on both the NarrativeReason and Timeline summarisation datasets. In this mixed dataset, we prepend a different prompt to clarify the task, ensuring that Phi can learn to apply its knowledge accordingly to each type of task. For details see appendix  A.3.
$\text{Phi}_{temp}$ and $\text{Phi}_{tl}$: We fine-tune Phi on the NarrativeReason and timeline summarisation datasets separately and obtain models $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$. We use these two models to generate timeline summaries for comparison, and examine whether fine-tuning on a single dataset can help the timeline summarisation task.
$\text{Phi}_{ICL}$: We use in context learning (ICL) to guide Phi to generate summaries. We provide the model with a pair consisting of a timeline and its corresponding summary as an example, and then let it generate summaries for other timelines.
LLaMA: We prompt LLaMA-3 (without any fine-tuning) to generate mental health related timeline summaries in a zero shot fashion.
TH-VAE: We use TH-VAE Song et al. (2024) to generate mental health summaries for comparison, while we use LLaMA-3 to generate high-level summaries.
@@ 4.3 Evaluation
We work with the timeline summaries from Song et al. (2024)4.1). Like Song et al. (2024), we employ Factual Consistency (FC), to measure how consistent timeline summaries are with the original timelines, and Evidence Appropriateness (EA), to measure the consistency between human written summaries and corresponding timeline summaries Song et al. (2024). Here, we use the annotated timeline evidence from Song et al. (2024) to directly generate high-level summaries. By contrast Song et al. (2024), generated clinically meaningful summaries from the timeline by first using TH-VAE to generate evidence related to mental health before generating high-level summaries. Since the goal of our paper is to validate the role of temporal reasoning in timeline summarisation, we directly use the annotated evidence to generate the high-level summary as the point is to make better use of the evidence rather than extract it from scratch.  [^2] We obtained the annotated dataset from the authors Song et al. (2024).
For human evaluation, we worked with two clinical psychology graduate students fluent in English to evaluate 30 summaries generated from 30 timelines (TalkLife). We follow the metrics used in Song et al. (2024) to evaluate the summaries from the perspectives of Factual Consistency and Usefulness (general/diagnosis/inter-&Interpersonal/MOC). A factually consistent summary should accurately represent the content of the timeline. We also evaluate the quality of a mental health summary on the basis of: General usefulness (contains the most clinically important information from the timeline in understanding a patient’s condition); Diagnosis (provides useful information about an individual’s mental state); inter-&Interpersonal (provides helpful information about an individuals’ needs and relationship patterns); MOC (provides useful information about an individual’s changes over time ).
@@ 5 Results and Discussion
Automatic evaluation: We conducted experiments with different combinations of KD methods and fine-tuning strategies (Table 1). Since we didn’t change output dimensions when fine-tuning LLaMA, CRD was not used during the KD process. Among individual methods, PKT performed the best, but combining PKT with NST achieved the best overall results. P-Phi and L-Phi are the best performing models with identical EA (correspondence to human summaries). L-Phi seems to have lower FC, which would suggest it is less faithful to the original timeline. However, human evaluation in Table 2 indicates that clinical psychologists have a clear preference for summaries generated by L-Phi, on all aspects. This suggests distilling information from a larger LLM is indeed beneficial.
@@
Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of the different models. Higher is better. Best results are in bold and second-best results in underline.
Regarding fine tuning strategies for timeline summarisation, the results for $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues. Notably, $\text{Phi}_{temp}$ performed the worst on both FC and EA metrics, suggesting that incorporating temporal reasoning information interferes with the LLM’s ability to effectively handle the timeline summarisation task. Additionally, in $\text{Phi}_{joint}$, combining the two types of data directly during training failed to integrate them effectively. As a result, the performance of $\text{Phi}_{joint}$ was worse than just using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
Human evaluation: Based on the results of the automatic evaluation, we selected the best-performing L-Phi and P-Phi models. Additionally, we included the non-fine-tuned zero-shot versions of LLaMA and Phi. This can help us understand in which specific aspects the model has improved with the inclusion of temporal reasoning information. The fine-tuned model L-Phi shows the greatest improvement in terms of factual consistency and usefulness (general). This aligns with our findings when analyzing the summaries, where the fine-tuned model significantly reduces hallucination, as shown in Table 2. In addition, we found that the fine-tuned model did not show significant improvement in terms of Moments of Change (MOC). In Appendix  A.4, we give examples of summaries generated by different models, these examples clearly demonstrate that the fine-tuned model can effectively reduce hallucinations.
Aspect Phi P-Phi LLaMA L-Phi Factual Consistency 2.90 3.32 3.58 3.83 Usefulness (General) 2.60 3.13 3.17 3.48    (Diagnosis) 2.90 3.37 3.45 3.62   —-(Inter-& Intrapersonal) 2.95 3.00 3.40 3.51   —-(MoC) 2.97 2.97 3.42 3.47
@@ 5.1 Why knowledge distillation works
[…]
To validate our hypothesis, we ran another set of experiments (JTRL) to analyze the internal representation difference between the two models. Our hypothesis here is that knowledge distillation results in better representations due to more polysemantic neurons, whose representations vary significantly to those of $\text{Phi}_{joint}$. To measure JTRL, we use the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) similarity score. CKA can be used to measure the similarity between internal representations of modelsConneau et al. (2020); Muller et al. (2021); Del and Fishel (2021); Moosa et al. (2023).
To calculate CKA, we used the same probing dataset. First, we calculate the sentence embeddings of each input by averaging the hidden state representation of the tokens. Then, we calculate the CKA similarity score between the mean sentence embeddings and each layer representation of the model. We did this both for the individual models and between models’ layers. From Figure 4, we can see that $\text{Phi}_{tl}$ shows a gradual increase in CKA across layers, peaking in the mid-to-late layers. This indicates that as the layers progress, $\text{Phi}_{tl}$ preserves and refines the information from the initial embeddings in a task-relevant way. The is likely due to KD encouraging this alignment by transferring task-relevant knowledge from the teacher. While $\text{Phi}_{joint}$ shows an initial increase in CKA, it saturates and flattens early. This shows failure to refine representations effectively in deeper layers, presumably due to conflicting objectives between the tasks. The lower CKA in later layers suggests that the model moves away from the initial embeddings in a less effective way for task-specific learning. Lastly, the CKA values between models show that they learn vastly different representations. In short, $\text{Phi}_{tl}$’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
@@
(b) UMAP projection $\text{Phi}_{tl}$
@@
Figure 4: CKA similarity score of both within and between $\text{Phi}_{tl}$ and $\text{Phi}_{joint}$ model representations.
@@ 6 Conclusions
We have created a dataset (NarrativeReason), containing relation triples to represent event sequences within narratives. We fine-tune a large LLM on NarrativeReason, then use knowledge distillation (KD) to transfer its enhanced temporal reasoning capability to a smaller LLM while leveraging the distilled knowledge to enhance performance in timeline summarisation tasks. We apply the model to a different domain from the one it was trained on, namely generating mental health related timeline summaries. Our results demonstrate that our KD approach produces more accurate summaries while significantly reducing hallucinations. Analysing the model’s internal representations shows that KD leads to better feature representations within the model, more aligned with timeline summarisation.
@@ A.2 Experiment on teacher model
After fine-tuning the teacher model, we used the TEMPREASON dataset tan2023towards to evaluate its performance pre- and post-fine-tuning. Specifically, we focused on the L3 part of the dataset, which deals with event-event relations, to determine whether the model could accurately infer the temporal sequence between two events. We use F1 as the evaluation metrics, the experimental results show that the fine-tuned model achieved a 0.07 improvement in this metric compared to the pre-fine-tuning model, highlighting its enhanced ability to infer event-event temporal relations.
A.3 Joint learning
When fine-tuning Phi using joint learning, we mix the timeline summarization and temporal reasoning datasets and shuffle their order. We then fine-tune Phi using this mixed dataset. For each data instance in this dataset, we prepend a prompt to clarify the task, ensuring that the LLM understands which task each data instance belongs to. For example, for timeline summarization data, we add the prompt "This is a timeline summarization task, your task is to summarize the provided timeline + timeline summarisation text". Similarly, for temporal reasoning tasks, we would prepend a different prompt to clarify the task at hand, ensuring that Phi can learn to apply its knowledge accordingly to each type of task.
A.4 Examples of summaries generated in different models
Gold (Diagnosis)This teenager girl had attempted suicide and is at high risk of trying again. She is depressed and lonely. She engages in self-hurt, like cutting herself, not eating, and abusing medication. She is being bullied at school, physically and sexually harassed. In addition, she defines herself as pansexual, (though she feels confused about this identity), which is not excepted at home by her mother. She also fears not being accepted by her friends. Furthermore, she has difficulty sleeping and suffers from nightmares. The home environment is stressful. Her parents argue about her situation, which triggers her stress and panic. She has bad experience with therapy and healthcare, and is suspicious towards the system. Her sense of agency is low. (IIP)She is longing for love and companionship. The other is perceived as unhelpful, uncaring, intrusive, and threatening. Her parents are especially seen as not understanding, unaccepting, controlling, and intimidating. The response of self is self-pity, self-blame, and hopelessness. (MOC)There is no sign of improvement in her mental health during the timeline, but she has been evaluated and will receive antidepressant medicine soon.
L-Phi (Diagnosis)The individual seems to be grappling with significant emotional turmoil, likely stemming from depression or anxiety disorders. They convey uncertainty about their sexual identity, coupled with recurring negativity from others, leading to challenges in maintaining social connections. Evidence of self-injurious actions such as self-mutilation and thoughts of ending one’s life have been observed. This person has faced relentless torment within educational settings, intensifying their alienation and despondency. Although their past interactions with mental health professionals aren’t directly mentioned, it can be inferred that they have sought assistance previously due to their consistent references to obtaining help. Fluctuations in their psychological well-being appear to be influenced by family disputes and insensitivity from acquaintances, further complicating their situation. (IIP)The individual grapples with depression, anxiety, and suicidal thoughts, manifesting feelings of despair, loneliness, and apprehension about being judged by loved ones and acquaintances. These experiences highlight their urgent need for empathy and comprehension from those around them, yet they find themselves subjected to criticism and misunderstanding owing to their sexual orientation and mental wellbes. Consequently, this person demonstrates limited self-forgiveness and has trouble acknowledging personal weaknesses and reaching out for assistance. (MOC)The overall mood and functioning of the individual remains consistently low throughout the timeline, with no clear evidence of any significant positive change. Additionally, there appears to be no specific event that triggers any noticeable change in the individual’s mood or well-being.
P-Phi (Diagnosis) Based on the provided social media posts, the individual appears to be struggling with several mental health challenges including depression, anxiety, self-harm tendencies, and possible substance misuse. The summary below provides a comprehensive overview of these issues as observed from the data available in the documents: The individual displays signs of depression through frequent expressions of sadness, hopelessness, and lack of motivation. Anxiety is evident in their constant worrying about future events and excessive fear responses. Self-harm tendencies are apparent in some of the social media posts where they mention engaging in harmful behaviors such as cutting or burning themselves. Lastly, there are indications of possible substance misuse, as they frequently post about using drugs or alcohol to cope with their emotional struggles. Overall, this person seems to be dealing with significant mental health concerns that require professional intervention and support. (IIP)The user wishes to find a partner who accepts them as they are while expecting people to understand and respect their sexuality. They seek support and understanding from those close to them but face challenges such as their family’s lack of understanding and acceptance of their sexuality, leading to feelings of isolation and longing for connection. This has resulted in the user experiensiong anxiety, depression, and self-harm tendencies, displaying signs of distress and disconnection within their immediate environment. However, despite these difficulties, the user demonstrates resilience and hope for finding acceptance and understanding in the future.", (MOC)The overall mood and functioning of the individual remains consistently low throughout the timeline, with no clear evidence of any significant positive change. Additionally, there appears to be no specific event that triggers any noticeable change in the person’s mood or well-being.
LLaMA (Diagnosis) This individual presents with severe depression, characterized by persistent suicidal ideation, feelings of worthlessness, and hopelessness. They also express intense loneliness, confusion about their sexual orientation, and distress from interpersonal conflicts at home and school. These factors contribute to their emotional pain, exacerbated by triggers related to romantic relationships and social acceptance. The individual has engaged in self-harm behaviors, including cutting, and experiences frequent suicidal ideation, indicating a high risk of suicide. They have shown some awareness of the need for professional help, having sought therapy and started anti-depressant medication, although they currently lack regular therapeutic support due to their therapist’s absence. Physical health concerns arise from neglecting nutrition and engaging in self-harming behaviors, while there is no specific mention of other physical health issues. The individual demonstrates low motivation for change and limited coping strategies, relying heavily on romantic relationships for validation and occasionally misusing prescription medication. Despite these challenges, it is crucial to address their immediate safety risks and facilitate access to comprehensive mental health care, including therapy, appropriate medication management, and support groups that can provide both emotional understanding and practical assistance in managing their symptoms and improving their quality of life.
(IIP)<<ANALYSIS>> The individual displays a deep-seated wish for a romantic partnership, frequently expressing a desire for a g̈irlfriend or boyfriend.T̈hey grapple with their sexual orientation, identifying as potentially pansexual but expressing uncertainty about their own understanding. Their pressing need for companionship and support is apparent, highlighted by their loneliness and fear of revealing their true self to friends. Moreover, they face considerable mental health challenges, including suicidal ideation, self-harming behaviors, and difficulty managing their prescribed medications. Interactions with peers are perceived negatively, with the individual reporting bullying and feeling victimized rather than supported. There is no evidence of recognizing complexity in others’ behavior; instead, social exchanges are colored by negativity and conflict. Their response to others is marked by fear, vulnerability, and a pervasive sense of being misunderstood. The fear of coming out to friends due to potential rejection or negative reactions underscores a profound lack of confidence in interpersonal relations and a heightened sensitivity to judgment from others. On a deeply personal level, the individual experiences overwhelming despair and self-blame. They consistently berate themselves (Ï’m such a screw-up)̈, burdened by guilt (Ït’s all my fault)̈, reflecting a severe erosion of self-esteem and self-worth. This self-criticism is coupled with a high risk of self-harm and persistent thoughts of suicide, demonstrating a pronounced inability to practice self-compassion and accept their own strengths and weaknesses. A dysfunctional intrapersonal pattern. (MOC)The individual’s emotional state fluctuates significantly between periods of profound despair and intermittent moments of hopefulness. These shifts are closely tied to various stressors, including family conflicts, school-related issues, and personal identity struggles.
Phi (Diagnosis)As an AI developed by Microsoft, I am programmed to follow strict ethical guidelines prioritizing user safety and privacy. Consequently, I can only analyze personal data in a secure and consensual environment specifically tailored for professional use. In a hypothetical scenario where such conditions are met, a trained professional would cautiously assess the situation, taking care to uphold confidentiality and abide by legal standards. This case involves an individual grappling with various mental health issues, including depression, anxiety, potential self-harm behaviors, and profound despair. Their struggle extends to uncertainties about sexual orientation, experiences of bullying, familial discord, and negative remarks stemming from disclosed private information, all contributing to significant distress over body image and self-worth. Alarmingly, they have openly expressed thoughts of suicide, underscoring the critical need for immediate intervention by certified mental health experts. It must be emphasized that this description serves as an example and does not substitute for actual therapeutic engagement with a licensed practitioner.
(IIP) Based on the analysis of the individual’s social media posts, several intrapersonal and interpersonal patterns have been identified that require careful consideration. The person has a profound need for connection and understanding about their sexual identity, frequently encountering dismissive reactions and misinterpretations from others which exacerbate their sense of confusion and loneliness. Their online communications reveal a struggle between a deep yearning for authentic connections and a pervasive fear of being judged, leading to fragmented expression of thoughts, especially when discussing sensitive issues like suicidal ideation and self-harm. This inner turmoil manifests externally through mixed signals expressing dark humor alongside serious statements indicative of self-directed hostility and despair. Yet, despite this volatility, they showcase resilience by acknowledging the necessity of medical assistance and therapy, suggesting a complex emotional landscape wherein adaptive behaviors coexist with maladaptive ones. Their response to external stimuli ranges from experiencing acute distress due to perceived betrayals and ridicule to demonstrating a proactive approach towards mental health management. It becomes apparent that while they grapple with severe depression, self-harm tendencies, and possible substance abuse, they simultaneously seek solace within specialized communities and display sporadic but meaningful strides toward recovery. Given the gravity of their situation, marked by life-thyesing thoughts and harmful coping mechanisms, urgent and multifaceted interventions are imperative. Such measures should encompass crisis stabilization, continuous psych. (MOC)The individual’s mood, well-being, and functioning exhibit significant fluctuations over time, transitioning between periods of despair and suicidal ideation as well as moments when they seek connection and support. These emotional shifts have been influenced by various events such as the loss of therapy access, familial conflicts, discovering potential partners with similar sexual orientation preferences, and a breach of privacy resulting from others reading their personal notes. However, there have also been positive developments including contemplating medication adjustments and identifying commonality in sexual orientation among their peer group, which has contributed to some degree of stability and improvement in their mental state. Despite these advancements, negative experiences persist, particularly marked by increased suicidal thoughts following the cessation of therapeutic sessions and heightened family discord. This complex interplay of factors underscores the importance of closely monitoring the patient’s evolving needs and responses to ongoing treatments.
Table 3: We have highlighted hallucinations and obvious errors in yellow. Compared to the gold summary, the summary generated by Phi still lacks some key information, especially in the ’Diagnosis’ part.